OpenStudy (anonymous):

Find the perimeter of a regular hexagon with sides of length 13.

OpenStudy (anonymous):

a hexagon has 6 sides so 6(13) = 78

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Find the area of an equilateral triangle with sides of length 8. in radical form

OpenStudy (anonymous):

 so Use Heron's Area Formula\[A = \sqrt{s(s-a)(s-b)(s-c)}\]\[s = (a+b+c)/2\]

OpenStudy (anonymous):

\[A = \sqrt{12(8)(8)(8)}\]\[A = 32\sqrt{6}\]